Assume the reserve requirement is 10% and no excess reserves are held. If an initial cash deposit of $10,000 is made, the money supply has the potential to increase by:

A.

$90,000.

B.

$110,000.

C.

$100,000.

D.

$11,000.
Answer #1

Multiplier = 1/r

= 1/0.1 = 10

The money supply would increase by:

= 10000*10

= 100000

option(C)